OpenStudy (whtcutie):

I need help with this question.... Factor the expression 3x^3+4x^2-15x-20. show work.

OpenStudy (decentnabeel):

\[3x^3+4x^2-15x-20\] \[=\left(3x^3+4x^2\right)+\left(-15x-20\right)\] \[\mathrm{Factor\:out\:}x^2\mathrm{\:from\:}3x^3+4x^2\mathrm{:\quad }x^2\left(3x+4\right)\] \[\mathrm{Factor\:out\:}-5\mathrm{\:from\:}-15x-20\mathrm{:\quad }-5\left(3x+4\right)\] \[=x^2\left(3x+4\right)-5\left(3x+4\right)\] \[\mathrm{Factor\:out\:common\:term\:}\left(3x+4\right)\] \[=\left(3x+4\right)\left(x^2-5\right)\] \[x^2-5=\left(x+\sqrt{5}\right)\left(x-\sqrt{5}\right)\] \[=\left(3x+4\right)\left(x+\sqrt{5}\right)\left(x-\sqrt{5}\right)\]

OpenStudy (mrs.ambrose614):

I am about to help you please wait before you hand out the medal plz

OpenStudy (mrs.ambrose614):

The awnser I came up with is -6x+2y-20 you have to simplify the expression

OpenStudy (triciaal):

we only have a few more weeks of Openstudy but when you answer questions please try and let the asker participate He/she is not even online according to the system.

OpenStudy (whtcutie):

thank you DecentNabeel but I don't really get it. Mrs.ambrose614, triciaal has a point I would like to be explain the presses and I would like to participate in order for me to understand the work. thank you all for trying to help.

jhonyy9 (jhonyy9):

@DecentNabeel great ,nice work really !!!

OpenStudy (whtcutie):

@3mar I need help with this question.

OpenStudy (3mar):

Well, I am here. Any Help... Any Time... How could I help you?

OpenStudy (whtcutie):

Factor the expression 3x^3+4x^2-15x-20 show work.

OpenStudy (3mar):

Any ideas? I think we will proceed with "grouping"!

OpenStudy (whtcutie):

so would it be (3x^3+4x^2)+(-15x-20)?

OpenStudy (3mar):

where this would lead you?

OpenStudy (whtcutie):

x^2(3x+4)-5(3x+4)?

OpenStudy (3mar):

then?

OpenStudy (3mar):

can you take (3x+4) as a common factor?

OpenStudy (whtcutie):